# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v7

@@ Abstract
[Objective] This study focuses on addressing the current lack of a unified formal theoretical framework in machine learning, as well as the deficiencies in interpretability and ethical safety assurance.
[Methods] A formal information model is first constructed, utilizing sets of well-formed formulas to explicitly define the ontological states and carrier mappings of typical components in machine learning. Learnable and processable predicates, along with learning and processing functions, are introduced to analyze the logical deduction and constraint rules of the causal chains within models.
[Results] A meta-framework for machine learning theory (MLT-MF) is established. Based on this framework, universal definitions for model interpretability and ethical safety are proposed. Furthermore, three key theorems are proved: the equivalence of model interpretability and information recoverability, the assurance of ethical safety, and the estimation of generalization error.
[Limitations] The current framework assumes ideal conditions with noiseless information-enabling mappings and primarily targets model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing the critical challenges currently faced in machine learning.
@@ 1 Introduction
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning [1, 2, 3, 4, 5]. Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
@@
Third, the mathematical foundations of interpretability tools are weak [11, 12, 13, 14, 15]. Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
@@
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learning[16, 17] . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systems[18, 19, 20, 21, 22, 23].
On this basis, this paper rigorously defines higher-order logical systems using recursive methods and establishes a formalized expression of states via the interpretation of sets of well-formed formulas, thus fully clarifying the mathematical essence of state representation in the definition of information. This fills the last theoretical gap in OIT and provides more powerful tools for applying OIT to analyze various problems in information science. Subsequently, the theories and methods of OIT are applied comprehensively—from prior knowledge encoding and data-driven training to model generalization applications—unifying core concepts such as prior distribution, data distribution, and parameter updates into a formalized state representation for the first time. This establishes a mathematical logic system that spans the entire machine learning process. Based on this, formal logical deduction and constraint rules for the causal chains in typical stages such as formal modeling and realization of training information, formal modeling and realization of model learning, and information processing and learning optimization in models are demonstrated, forming the Meta-Framework for Machine Learning Theory (MLT-MF).
Finally, MLT-MF is applied to several hot topics in machine learning, establishing universal definitions for model interpretability and ethical safety, and proving three important theorems: the equivalence of model interpretability and information recoverability, ethical safety assurance, and generalization error estimation. This provides a powerful tool based on a unified formal logical framework for both theoretical analysis and engineering practice in machine learning.
@@ 2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability. In response,[22] proposes four Postulates of information:
@@
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realization[24, 25, 26].
@@ 2.2 Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more concise[27]. Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 2 (Interpretation of a Formal System): An interpretation $E$ of formal system $\mathcal{L}$ over a domain $D_{E}$ consists of a non-empty set $D_{E}$, a specified set of elements $(\widehat{a}_{1},\widehat{a}_{2},\ldots)$ over $D_{E}$, a set of functions $(\widehat{f}_{i}^{n},n>0,i>0)$ over $D_{E}$, and a set of relations $(\widehat{A}_{i}^{n},n>0,i>0)$ over $D_{E}$ and the specified elements.
According to this definition, the variables $x_{1},x_{2},\ldots$ in $\mathcal{L}$ are interpreted as objects. The set $D_{E}$ is the domain over which variables take their values. The set of specfied elements corresponds to the special objects denoted by the individual constants in $\mathcal{L}$. Functions and relations provide concrete interpretations for the function and predicate symbols in $\mathcal{L}$.
@@
(1) Definitional Invariance: The definitions of the object set and the time set remain invariant over the entire domain.
(2) Expressivity of Attributes: All attributes of the object set—including properties, structures, values, and relations—over the domain can be expressed via functions and predicates within a formal system conforming to Definition 1.
(3) Expressivity of Rules: The evolution rules governing various attributes of the object set can be expressed, without implicit self-reference, via temporal variables and application connectives in a formal system as specified in Definition 1.
(4) Logical Consistency: There exists no logical contradiction among the various attributes and evolution rules of the object set, nor among any possible logical conjunctions thereof.
@@
Theorem 1 (Formal Expression of States): Let $S(X,T)$ be the state set of an object collection $X$ over a specified time set $T$, satisfying Postulate 5. Then $S(X,T)$ can always be represented as the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$, and is logically consistent.
Proof: Since $S(X,T)$ satisfies Postulate 5, we have:
  • By (1), the elements of $X\times T$ are well-defined, so $S(X,T)$ has a definite domain.
@@
  • By (2), all attributes of $X$ over $T$ in $S(X,T)$ can be expressed by functions and predicates in a formal system $\mathcal{L}$ as specified in Definition 1.
  • By (3), suppose $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ and $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$ are two attributes in $S(X,T)$ pertaining to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. If every $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ evolves from $t_{1}$ to $t_{2}$ according to some rule into $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$,
    then $(\forall A_{1})(\forall f_{1})(\forall x_{1})(A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))\rightarrow\exists A_{2}\exists f_{2}\exists x_{2}(A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))))$ is a well-formed formula in $\mathcal{L}$ that, as an interpretation over $x_{1},x_{2},t_{1},t_{2}$, expresses the corresponding rule in $S(X,T)$. Other rules in $S(X,T)$ can likewise be expressed through corresponding well-formed formulas in $\mathcal{L}$ interpreted over $X\times T$.
  • Let the set of such well-formed formulas in $\mathcal{L}$ representing $S(X,T)$ be defined as its interpretation ${E}$ over the domain $X\times T$.
  • By (4), no logical contradiction exists among any attributes or rules in $S(X,T)$, nor among any logical conjunctions thereof. Thus, as a representation of $S(X,T)$, both ${E}$ and its logical closure are logically consistent.
@@
In summary, $S(X,T)$ is indeed an interpretation ${E}$ of a set of well-formed formulas in $\mathcal{L}$ over the domain $X\times T$, and both itself and its logical closure are logically consistent. The theorem is thus proved.
@@
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses [28] further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
@@ 2.3 Finite Automaton States
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n_{T}\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory [29], $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n_{T}-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n_{r}\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n_{u}\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
@@
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$.
@@
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$.
@@
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$.
@@
$S(M,T)=Cn(\{\varphi_{U}(t_{i})\}\cup\{\varphi_{I}(t_{i})\}\cup\{\varphi_{O}(t_{i})\}\cup\{\varphi_{\delta}(t_{i})\}\cup\{\varphi_{\lambda}(t_{i})\})$ (2.6)
@@ 2.4 Definition of Information
As shown in[22], the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
@@
Definition 4 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, that is,
$I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ (2.7)
or
$I\left(S_{o}(o,T_{h})\right)\underline{\triangleright}S_{c}(c,T_{m})$ (2.8)
The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
[…]
This shows that the carrier state set $S_{c}(c,T_{m})$ is always induced by the ontological state set $S_{o}(o,T_{h})$, and thus the enabling mapping $I$ must be surjective. For simplicity, unless otherwise noted, subsequent discussions assume that $I$ is a single-valued mapping.
Therefore, for any $\psi\in S_{c}(c,T_{m})$, there exists at least one $\varphi\in S_{o}(o,T_{h})$ such that $I(\varphi)\underline{\triangleright}\psi$. In this case, we write $I^{-1}(\psi)\underline{\triangleleft}\varphi$ or $\varphi\underline{\triangleright}I^{-1}(\psi)$ to preserve the consistency of the enabling direction.
@@
Reference [24]is generated as a foundational work in information theory. The paper proposed that "$H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $01\leq i\leq n,\tau>$ 0).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,

### v9

@@ Abstract
[Objective] This study addresses key challenges in machine learning, namely the absence of a unified formal theoretical framework and the lack of foundational theories for model interpretability and ethical safety.
[Methods] We first construct a formal information model, explicitly defining the ontological states and carrier mappings of typical machine learning stages using sets of well-formed formulas. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the causal chain logic and constraint laws governing machine learning processes.
[Results] We establish the Machine Learning Theory Meta-Framework (MLT-MF), on which we further propose universal definitions for model interpretability and ethical safety. We prove and validate three key theorems: the relationship between model interpretability and information existence, ethical safety assurance, and the upper bound estimation of total variation distance (TVD).
[Limitations] The current framework assumes ideal, noise-free information enabling mappings and focuses primarily on model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing critical issues in contemporary machine learning.
@@ 1 Introduction
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning Rob2021 ; Marius2023 ; Meng2023 ; Alquier2024 ; Vilalta2002 . Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
@@
Third, the mathematical foundations of interpretability tools are weak Ribeiro2016 ; Lundberg2017 ; Venkata2025 ; Lee2023 ; Alvarez2018 . Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
@@
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learningShwartz2017 ; Xu2022 . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systemsXu2014 ; xu2015 ; Jianfeng2021 ; Jianfeng2023 ; Xu2024 ; Xu2025 .
On this basis, we rigorously define higher-order logic systems using recursion, and establish formal state expressions through interpretations of sets of well-formed formulas (WFFs), completing the last theoretical gap in OIT regarding the mathematical essence of state expressions. We then thoroughly apply the theory and methods of OIT to unify the core concepts throughout the machine learning lifecycle—prior knowledge encoding, data-driven training, and model generalization—as formal state expressions; we establish a mathematical logic system covering the entire ML process. Using this, we analyze and prove the causal-chain logic and constraint rules for typical stages in machine learning, including the formal modeling and realization of training information, the formal modeling and realization of model learning, and the model’s information processing and learning optimization, thus forming a Meta-Framework for Machine Learning Theory (MLT-MF). Finally, using MLT-MF, we propose universal definitions of model interpretability and ethical safety, prove and validate—via a simple feedforward neural network—three important theorems: model interpretability and information existence, ethical safety assurance, and an upper bound estimate on total variation distance (TVD). These results provide a powerful tool, based on a unified formal logic framework, for both theoretical analysis and engineering practice in machine learning.
@@ 2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability.
To address this, we first propose the concept of an enabling mapping.
Definition 1 (Enabling Mapping) Let $X$ and $Y$ be two sets, and let $f$ be a mapping. If for any $x\in X$, there exists a $y\in Y$, and for any $y\in Y$, there exists at least one $x\in X$ such that $y=f(x)$, and furthermore, the existence of $y$ is physically realized only because of $x$, then $f$ is called an enabling mapping from $X$ to $Y$. This is denoted as:
$f:X\rightrightarrows Y$ (2.1)
or
$f(X)\underline{\triangleright}Y$ (2.2)
[…]
If $f(X)\underline{\triangleright}Y$, we also denote this as $f^{-1}(y)\underline{\triangleleft}x$ or $x\underline{\triangleright}f^{-1}(y)$, to maintain consistency in the directionality of enabling.
Based on the concept of enabling mapping, Xu2024 proposed four postulates of information:
@@
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realizationShannon1948 ; Landauer1996 ; Wiener2019 .
@@ 2.2 Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more conciseHamilton1988 . Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 3 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
@@
(1) Domain invariance: definitions of the object set and time set remain unchanged.
(2) Numerical determinacy: concrete numerical values determine values of certain properties and forms.
(3)Attribute expressibility: properties, forms, values, relations are expressible via functions and predicates in a formal system satisfying Definition 2.
(4) Rule expressibility: evolution rules among attributes are expressible in such a formal system using time variables and implication, and contain no implicit self-reference.
(5) Logical coherence: no logical contradictions exist among attributes and rules, nor among any logical conjunctions thereof.
@@
Theorem 1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate 5, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate 5, we construct the corresponding formal system $\mathcal{L}$ as follows:
By (1), the definitions of the elements in $X$ and $T$ remain unchanged, ensuring determinacy. The variables in the formal system $\mathcal{L}$ are defined to correspond to each element of $X\times T$, thus making $X\times T$ the domain of discourse for $\mathcal{L}$. Consequently, $S(X,T)$ has a well-defined domain.
By (2), the constants in $\mathcal{L}$ are defined to represent all possible values in $S(X,T)$.
@@
By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$. These properties can be expressed through functions and predicates in the formal system $\mathcal{L}$ that conform to Definition 1.
By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.And then let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$. The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))\rightarrow\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2})))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$. For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
@@
Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The theorem is thus proved.
Note that Theorem 1 stipulates that variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$. In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
@@
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses qiu2025 further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
@@ 2.3 Finite Automaton States
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory mathsea2002 , $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
@@
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$$i=1,\ldots,n$.
@@
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$$i=1,\ldots,n-1$.
@@
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$$i=2,\ldots,n$.
@@
$S(M,T)=$ $\{\varphi_{ij}(t_{i}),\varphi_{I}(t_{j}),\varphi_{O}(t_{k}),\varphi_{S}(t_{j}),\varphi_{A}(t_{j}),$ (2.8)
$i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}$
@@ 2.4 Definition of Information
As shown inXu2024 , the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
@@
Definition 5 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$.The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
@@
Reference Shannon1948 is generated as a foundational work in information theory. The paper proposed that $H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $1\leq i\leq n,\tau>0$).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,
@@
2.5 Transmissibility of Information
Xu2014 ; xu2015 discuss transmissibility using composite mappings. Building on Definitions 1 and 5, we can characterize the fundamentals of information transmission more rigorously:
Theorem 3 (Transmissibility of Information) If $I_{1}=\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\rangle$ is information from an ontology $o$ to a carrier $c_{1}$, and $I_{2}=\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is information from carrier $c_{1}$ to carrier $c_{2}$, then the composition $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is also information, which we call transmission of information from $I_{1}$ via carrier $c_{1}$ to carrier $c_{2}$.
Proof: According to Definition 1, Definition 5, and the given conditions, for any $S_{o}(x,t)\in S_{o}(o,T_{h})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Furthermore, there exists $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$. We define $I_{3}(S_{o}(x,t))=S_{c_{2}}(x_{2},t_{2})$. Since the physical realization of $S_{c_{2}}(x_{2},t_{2})$ depends on the existence of $S_{c_{1}}(x_{1},t_{1})$, and the existence of $S_{c_{1}}(x_{1},t_{1})$ in turn depends on $S_{o}(x,t)$, it follows by transitivity of causality that the realization of $S_{c_{2}}(x_{2},t_{2})$ ultimately depends on $S_{o}(x,t)$.
Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Thus, $I_{3}(S_{o}(x,t))=I_{2}(I_{1}(S_{o}(x,t)))\underline{\triangleright}I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
Therefore, $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 5. In essence, this is the information $I_{1}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$. The theorem is thus proved.
Transmission is an important way of information flow. Communication, processing, input, and output are all concrete manifestations of information transmission. In many cases, information undergoes more complex flows through chain-like transmission mechanisms. From Theorem 3, we can readily derive the following:
Corollary 1 (Information Chain) Let $\{I_{i}=\langle o,T_{h},S_{o},c_{i},T_{m_{i}},S_{c_{i}}\rangle|i=1,\ldots,n\}$ be a sequence of information mappings, where for any $i>1$, it holds that $c_{i-1}=o$, $T_{h}=T_{m_{i-1}}$, and $S_{o}(o,T_{h})=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any $1\leq i\leq j\leq n$, $I_{ij}=\langle o,T_{h},S_{o},c_{j},T_{m_{j}},S_{c_{j}}\rangle$ also qualifies as information according to Definition 5.